OpenStudy (anonymous):

Helen is helping plan a trip for three classes at her son's school . School rules say that classes on trips must have1 adult for every 6 children one class has 24 children; one has 26 children; and the thrid has 28 children. Each classhas 1 teacher . How many OTHER adults are needed for the trip

OpenStudy (ash2326):

School rules say that classes on trips must have 1 adult for every 6 children. There are =24+26+28=78 children So we need \(\large \frac{78}{6}=13\) adults Already there are 1 teacher for each class= so 3 teachers are there Other adults needed=13-3=10 so 10 more adults are needed